This sequence encodes a protein that you wish to study further by cloning the gene. The protein of interest has a molecular weight of at least 10 kDa. The sequence was sequenced by a primer walking approach, but unfortunately the correct order of the sequence traces was lost (really bad lab notebook skills).

Trace 1

aacgagttaaggagccagcgtaccttcgcaccgccatacatgaattttcttggctttttctatgtggatggcaatagtctagagtcggacctgcaggcatgcaag

Trace 2

cagcctgttagtaggtcttactgagtcgggcgccgaattcgagctcggtacccggggatccatgagtgggcgccagttattcgtactattgggaggtcc

Trace 3

ccgggtattttgtattcaatattgaataaggaatttttcatgcagagaaaaggatgtttacggctcgagcgcactcgcacatataactgtcggcagaaacgagttaaggagc

Trace 4

tactattgggaggtccaaatggttttacgggagttgcactgggcgaatgctggagctattacgattcgtttaacatttccatatcgaattctcagacgactccgaatccgggtatttt

Trace 5

cctgcaggcatgcaagcttggcataggtcagttcatccagggtgatgggtgtatcgtttcaatgacccgattcggaacg

Answer the following--

Determine the correct order of the sequence traces and the entire sequence.

Locate an open reading frame that could encode the protein of interest.

Determine the amino acid sequence, the molecular weight, and the isoelectric point of this protein.

Design a set of PCR primers that will amplify at a minimum this entire coding region. These primers must pass all the requirements for good primers.

Develop a cloning strategy that would enable you to clone at a minimum the coding region from this sequence into the vector pUC19, so that the coding region is in a clockwise orientation. There are several ways to accomplish this cloning objective; the only requirement is that it is theoretically possible and that you can show that the coding region is in a clockwise orientation.
Answer #1

1. The trace sequences can be aligned together by pasting the entre set in a word document and searching the initial string of nucleotides of one trace sequence that would match with any of the other trace sequences. This way, The beginning and the end of each sequence would be known and the entire sequence can be merged. Once we do that, we find that the sequences should be in the following order: Trace 2 > Trace 4 > Trace 3 > Trace 1 > Trace 5. The matching sequences are highlighted in the same color in the image below:

image?w=602&h=444&rev=3&ac=1&parent=1i86

2. The ORF for the merged entire sesquence can be identified using the software ORF finder online tool in the NCBI website. The results would indicate differerent possible ORFs chosen at different frames and also the corresponding amino acid sequence and the number of amino acids. Since the size of the protein is given as ~10 KDa, and the approximate molecular weight of an amino acid is 110 Daltons, we need to look for an ORF that is at least 10000 / 110 = 91 amino acids. The first ORF displayed is 100 amino acids and it starts from the nucleotide 62 and ends at 364. The ORF with start and stop codons is given below:

image?w=602&h=144&rev=7&ac=1&parent=1i86

The corresponding amino acid sequence is :

>lcl|ORF1
MSGRQLFVLLGGPNGFTGVALGECWSYYDSFNISISNSQTTPNPGILYSI
LNKEFFMQRKGCLRLERTRTYNCRQKRVKEPAYLRTAIHEFSWLFLCGWQ

The molecular weight and isoelectric point of the above amino acid sequence can be determined using the online tool 'protparam' available in the ExPaSy website.

Protparam analysis of the above sequence has computed the molecular weight and pI of the protein to be

Molecular weight: 11609.37 Theoretical pI: 9.46

In order to design primers for the sequence, we would choose the region near the start codon to design the forward primer and the region near the stop codon to design the reverse primer. The primers should flank the entire ORF region.

>entire_sequence

Cagcctgttagtaggtcttactgagtcgggcgccgaattcgagctcggtacccggggatccatgagtgggcgccagttattcgtactattgggaggtccaaatggttttacgggagttgcactgggcgaatgctggagctattacgattcgtttaacatttccatatcgaattctcagacgactccgaatccgggtattttgtattcaatattgaataaggaatttttcatgcagagaaaaggatgtttacggctcgagcgcactcgcacatataactgtcggcagaaacgagttaaggagccagcgtaccttcgcaccgccatacatgaattttcttggctttttctatgtggatggcaatagtctagagtcggacctgcaggcatgcaagcttggcataggtcagttcatccagggtgatgggtgtatcgtttcaatgacccgattcggaacg

The primers should have more or less the same Tm. It is better for annealing if the 3' end of the primers have C or G nucleotides. Also, primers should have a GC content of about 40 to 60%. Tm or the melting temperature is calculated using the formula:

Tm = 2 * ( A + T) + 4 * (G + C)

Ideally, the Tm should be between 60\degreeC and 70\degreeC

Forward primer: 5'- ATGAGTGGGCGCCAGTTATTCG -3'

Tm = 2* (10) + 4*(12) = 20 + 48 = 68\degreeC

Reverse primer: 5'- CTAGACTATTGCCATCCACATAG -3'

Tm = 2* (13) + 4* (10) = 26 + 40 = 66\degreeC

Cloning:

Cloning of the above sequence can be done by identifying the right restriction enzymes. The restrction enzyme should not cut the ORF region, and for the integration of the DNA insert into the vector in the correct 5'>3' orientation, one should use directional cloning, where two restriction enzymes are used instead of one..

The first step to plan for cloning of the DNA is to identify the restriction sites in the MCS region of the given plasmid and analysing the above DNA sequence using the web tool, "webcutter".

Using the pUC19 vector map, we can identify the sites in the MCS region. The information can be obtained from NEB website. There are several restriction sites given but we could consider the commonly available enzyme sites. The (common) restriction sites in the MCS in 5' > 3' orientation are : Hind III > Sph1 > Pst1 > Sal1 > Xba1 > BamH1 > Xma1 > Sma1 > Kpn1 > Sac1 > Eco R1.

With Webcutter analysis, the list of restriction enzymes (that match with the MCS of pUC19) that cut the above sequence "only once" and those that do not cut the sequnence and their corresponding nucleotide positions are given below:

Hind III - 396

Sph 1 - 394

Pst1 - 388

Sal1- does not cut the sequence

Xba1- 371

BamH1- 62

Xma1- 57

Sma1- 59

Kpn1- 57

Sac1- 51

EcoR1- cuts twice, once once into the ORF

Now, we cannot use HindIII here as it would result in the DNA to integrate in the reverse orientation. The right set of enzymes to be used are Sal 1 and Xba1, Sal1 site can be inroduced in the forward primer and Xba 1 in the reverse primer, and the sequence can be amplified and cloned into pUC19.